OpenStudy (anonymous):

Find the slope and y-intercept of the line and draw its graph. 4x + 5y − 1 = 0 x,y intercept?

OpenStudy (anonymous):

slope =-4/5 y intercept= 1/5

OpenStudy (anonymous):

for the slope rearrange the equation to the form y=mx+b m is the slope, b is the y intercept\[5y=-4x+1\]\[y=\frac{-4}{5}x+\frac{1}{5}\]so slope =-4/5 y intercept=1/5

OpenStudy (anonymous):

x intercept=?

OpenStudy (anonymous):

x intercept is when y=0 so u substitute y=0 and solve for x

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

4x+5(0)-1=0 4x=1 x=1/4